# Principal stratification with $U$-statistics under principal ignorability

### v4

@@ 3.2 Estimators based on nonparametric identification formulas
Upon closer investigation, $\widehat{\tau}_{h,\pi+p}^{10}$ is based on the estimated treatment propensity score and the principal score, $\widehat{\tau}_{h,\pi+\mu}^{10}$ is based on the estimated treatment propensity score and the pairwise outcome mean, and $\widehat{\tau}_{h,p+\mu}^{10}$ is based on the estimated principal score and the pairwise outcome mean. For brevity, we will use $\tau_{h}^{10}$ as the running example for exposition in the remainder of the article. Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in the middle of Section S2 of the Online Supplement. These estimators have strong requirements on model specification to consistently estimate the PGCEs. To reiterate, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment propensity and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
Remark 2
It is worth noting that our weighting estimator $\widehat{\tau}_{h,\pi+p}^{s}$ leverages the inverse propensity score and the principal score for pairs, and can be viewed as an extension of the pairwise inverse propensity score weighted estimator in Mao (2017) for estimating $\tau_{h}$ in the absence of the intermediate variable. The theoretical basis for the consistency and asymptotic normality of $\widehat{\tau}_{h,\pi+p}^{s}$, $\widehat{\tau}_{h,\pi+\mu}^{s}$, and $\widehat{\tau}_{h,p+\mu}^{s}$, can be established by first recognizing their ratio structures and then investigating the numerators and the denominators separately; See the Online Supplement for additional details. The asymptotic behavior for the denominators involving $\widehat{p}_{z}$ can be obtained using empirical process theory for the parametric class (van der Vaart, 1998, §19.2). The numerators are parameter-indexed $U$-processes whose asymptotic properties can be established following Lemma S1 and Proposition 1 of Mao (2017).
@@ 4.1 The efficient influence function for the PGCE estimand
We first develop more efficient estimators with $\sqrt{n}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification. Second, the lower bound of the asymptotic variance is the variance of the EIF, providing a benchmark to define efficient estimators. Third, the EIF provides a device for incorporating data-adaptive machine learners for the nuisance parameters, suggesting feasible estimators that achieve the asymptotic variance lower bound with $\sqrt{n}$ convergence rate (Chernozhukov et al., 2018). We explore these properties by first formalizing the EIF for each PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|x)$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},X_{i})=(z,d,x)$, and $F_{X}(\cdot)$ denote the marginal distribution of $X$. We can further define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(x)h(Y,y)\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(x)h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|x)\allowbreak\mathrm{d}F_{X}(x)$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following Theorem 1 is obtained.
@@
$\phi_{h}^{10}(\mathcal{O})=$ $~\frac{ZDe_{10}(X)}{\pi(X)p_{1}(X)}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(X)\right\}+\frac{(1-Z)(1-D)e_{10}(X)}{\{1-\pi(X)\}\{1-p_{0}(X)\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(X)\right\}$
$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(X)+m_{0}^{10}(X)\right\},$
with $m_{1}^{10}(X)=\int e_{10}(x)\mu_{1100}(X,x)\mathrm{d}F_{X}(x)$, $m_{0}^{10}(X)=\int e_{10}(x)\mu_{1100}(x,X)\mathrm{d}F_{X}(x)$, and
@@
The EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1. The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al. (2022) with respect to the distribution of observed data, $\mathcal{O}$. Based on Theorems 1 and S3, the asymptotic variances of estimators for the PGCEs are lower bounded by the respective semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$ for $s\in\{10,00,11\}$.
@@ 4.2 Multiply robust estimators
The EIFs imply the semiparametric efficiency lower bound and pave the way for formulating optimal estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and require care for constructing the final estimator. We use $U$-statistics in constructing the final estimator, which could be shown to have the same influence function given in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that $\tau_{h}^{10}$ can be rewritten as
@@
$=2^{-1}\bigg[\frac{Z_{i}D_{i}e_{10}(X_{i};\beta)}{\pi(X_{i};\alpha)p_{1}(X_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(X_{j};\beta)}{\{1-\pi(X_{j};\alpha)\}\{1-p_{0}(X_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(X_{i},X_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(X_{j};\beta)}{\pi(X_{j};\alpha)p_{1}(X_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(X_{i};\beta)}{\{1-\pi(X_{i};\alpha)\}\{1-p_{0}(X_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(X_{j},X_{i};\gamma)\right\}$
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(X_{i},X_{j};\gamma)$
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(X_{j},X_{i};\gamma)\bigg],$
@@
which is essentially a $U$-statistic but in the form of ratios. Estimators $\widehat{\tau}_{h,\mathrm{tr}}^{00}$ and $\widehat{\tau}_{h,\mathrm{tr}}^{11}$ are similarly constructed and defined in Section S4.1 of the Online Supplement. The asymptotic properties of these estimators are summarized in Theorem 2.
@@
Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(X;\alpha^{*}),\pi(X;\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(X;\beta^{*}),\allowbreak p_{1}(X;\widehat{\beta}),1-p_{0}(X;\beta^{*}),1-p_{0}(X;\widehat{\beta})\}>\varepsilon$ for all $X\in\mathcal{X}$. Then, each estimator $\widehat{\tau}_{h,\mathrm{tr}}^{s}$, $s\in\{10,00,11\}$, is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
Under Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. This triple robustness property of $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is obtained by separately investigating the properties of its numerator and the denominator. Specifically, the denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via a nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
We propose further improvement of the parametric multiply robust estimator through debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov et al., 2018) or EIFs, but allows efficient estimation of causal effects with data-adaptive nuisance estimators. Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates. Compared to parametric working models, leveraging machine learning-based working models for the nuisance functions is more advantageous in terms of robustness against potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting—is adopted. In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov et al., 2018). Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
We adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2022). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq n\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al. (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. We use the following example to demonstrate the definition of $\mathcal{I}$ and $\mathcal{P}$.
@@
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we can obtain the following debiased machine learning estimator for $\tau_{h}^{10}$,
@@
The debiased machine learning estimators $\widehat{\tau}_{h,\mathrm{dml}}^{00}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ can be constructed in a similar fashion and are given in Section S5 of the Online Supplement. We now discuss the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$. We first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following Theorem 3 summarizes the convergence and efficiency properties of the debiased machine learning estimators.
@@
    $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
$\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al. (2022, Lemma 15) and Escanciano and Terschuur (2022, Lemma 4). Regulatory conditions in Theorem 3 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt and Dukes, 2022). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via nonparametric bootstrap.

### v5

@@ 3.2 Estimators based on nonparametric identification formulas
For brevity, we will use $\tau_{h}^{10}$ as a focused example in the remainder of the article. Results for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are analogous and are given in Section S2 of the Online Supplement. These estimators have strong requirements on model specification to estimate the PGCEs consistently. To reiterate, $\widehat{\tau}_{h,\pi+p}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+p}$, that is, when the treatment propensity and the principal score are both correctly specified; $\widehat{\tau}_{h,\pi+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{\pi+\mu}$, since the doubly robust estimator for $p_{z}$ is consistent under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}$; and $\widehat{\tau}_{h,p+\mu}^{s}$ for $s\in\{10,00,11\}$ are consistent under $\mathcal{M}_{p+\mu}$.
@@ 4.1 The efficient influence function for the PGCE estimand
We first develop more efficient estimators with $n^{1/2}$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickle1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF integrates multiple nuisance models and facilitates estimators that are robust against partial model misspecification. Second, the lower bound on the asymptotic variance is the variance of the EIF, providing a benchmark for defining efficient estimators. Third, the EIF provides a device for incorporating data-adaptive machine learners for nuisance parameters, yielding optimal estimators that achieve the asymptotic variance lower bound with a $n^{1/2}$ convergence rate (Chernozhukov2018). We explore these properties by first formalizing the EIF for each PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(\bm{X}_{1})e_{10}(\bm{X}_{2})\allowbreak\mu_{1100}(\bm{X}_{1},\bm{X}_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
where $\varphi_{h,N}^{s}(\mathcal{O})$ and $\varphi_{D}^{s}(\mathcal{O})$ are EIFs for $\tau_{h,N}^{s}$ and $\tau_{D}^{s}$, respectively. Let $F_{zd}(\cdot|\bm{x})$ denote the conditional distribution of $Y_{i}(z)$ given $(Z_{i},D_{i},\bm{X}_{i})=(z,d,\bm{x})$, and $F_{\bm{X}}(\cdot)$ denote the marginal distribution of $\bm{X}$. We define marginalized functions of the observed outcome, $\nu_{1,zd}^{s}(Y)=\int e_{s}(\bm{x})h(Y,y)\mathrm{d}F_{zd}(y|\bm{x})\allowbreak\mathrm{d}F_{\bm{X}}(\bm{x})$ and $\nu_{0,zd}^{s}(Y)=\int e_{s}(\bm{x})h(y,Y)\allowbreak\mathrm{d}F_{zd}(y|\bm{x})\allowbreak\mathrm{d}F_{\bm{X}}(\bm{x})$, for $z,d\in\{0,1\}$ and $s\in\{10,00,11\}$. Then, the following Theorem 1 is obtained.
@@
$\phi_{h}^{10}(\mathcal{O})=$ $~\frac{ZDe_{10}(\bm{X})}{\pi(\bm{X})p_{1}(\bm{X})}\left\{\nu_{1,00}^{10}(Y)-m_{1}^{10}(\bm{X})\right\}+\frac{(1-Z)(1-D)e_{10}(\bm{X})}{\{1-\pi(\bm{X})\}\{1-p_{0}(\bm{X})\}}\left\{\nu_{0,11}^{10}(Y)-m_{0}^{10}(\bm{X})\right\}$
$+\{\psi_{D_{1}}(\mathcal{O})-\psi_{D_{0}}(\mathcal{O})\}\left\{m_{1}^{10}(\bm{X})+m_{0}^{10}(\bm{X})\right\},$
with $m_{1}^{10}(\bm{X})=\int e_{10}(\bm{x})\mu_{1100}(\bm{X},\bm{x})\mathrm{d}F_{\bm{X}}(\bm{x})$, $m_{0}^{10}(\bm{X})=\int e_{10}(\bm{x})\mu_{1100}(\bm{x},\bm{X})\mathrm{d}F_{\bm{X}}(\bm{x})$, and
@@
For brevity, the EIFs for $\tau_{h}^{00}$ and $\tau_{h}^{11}$ are given in Theorem S3 of Online Supplement Section S3.1. The EIFs for the PGCEs are derived following the chain rule via the method of Gâteaux derivatives with respect to the distribution of observed data, $\mathcal{O}$ (Ichimura2022; Hines2022). Based on Theorems 1 and S3, the asymptotic variances of estimators for the PGCEs are lower bounded by the respective semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$ for $s\in\{10,00,11\}$.
@@ 4.2 Multiply robust estimators
The EIFs imply the semiparametric efficiency lower bound and pave the way for the formulation of optimal estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and require care for constructing the final estimator. We use U-statistics to construct the final estimator, which has the same influence function as in Theorem 1, thereby achieving local semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that $\tau_{h}^{10}$ can be rewritten as
@@
$=2^{-1}\bigg[\frac{Z_{i}D_{i}e_{10}(\bm{X}_{i};\beta)}{\pi(\bm{X}_{i};\alpha)p_{1}(\bm{X}_{i};\beta)}\times\frac{(1-Z_{j})(1-D_{j})e_{10}(\bm{X}_{j};\beta)}{\{1-\pi(\bm{X}_{j};\alpha)\}\{1-p_{0}(\bm{X}_{j};\beta)\}}\left\{h(Y_{i},Y_{j})-\mu_{1100}(\bm{X}_{i},\bm{X}_{j};\gamma)\right\}$
$\qquad\qquad+\frac{Z_{j}D_{j}e_{10}(\bm{X}_{j};\beta)}{\pi(\bm{X}_{j};\alpha)p_{1}(\bm{X}_{j};\beta)}\times\frac{(1-Z_{i})(1-D_{i})e_{10}(\bm{X}_{i};\beta)}{\{1-\pi(\bm{X}_{i};\alpha)\}\{1-p_{0}(\bm{X}_{i};\beta)\}}\left\{h(Y_{j},Y_{i})-\mu_{1100}(\bm{X}_{j},\bm{X}_{i};\gamma)\right\}$
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\mu_{1100}(\bm{X}_{i},\bm{X}_{j};\gamma)$
$\qquad\qquad+\{\psi_{D_{1}}(\mathcal{O}_{j};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{j};\alpha,\beta)\}\{\psi_{D_{1}}(\mathcal{O}_{i};\alpha,\beta)-\psi_{D_{0}}(\mathcal{O}_{i};\alpha,\beta)\}\mu_{1100}(\bm{X}_{j},\bm{X}_{i};\gamma)\bigg],$
@@
which is essentially a scaled U-statistic. Estimators $\widehat{\tau}_{h,\mathrm{tr}}^{00}$ and $\widehat{\tau}_{h,\mathrm{tr}}^{11}$ are similarly constructed and defined in Section S4.1 of the Online Supplement. The asymptotic properties of these estimators are summarized in Theorem 2.
@@
Under Assumptions 1 - 3, and suppose there exists some $\varepsilon\in(0,1)$ such that $\varepsilon<\{\pi(\bm{X};\alpha^{*}),\pi(\bm{X};\widehat{\alpha})\}<1-\varepsilon$, and $\{p_{1}(\bm{X};\beta^{*}),\allowbreak p_{1}(\bm{X};\widehat{\beta}),1-p_{0}(\bm{X};\beta^{*}),1-p_{0}(\bm{X};\widehat{\beta})\}>\varepsilon$ for all $\bm{X}\in\bm{\mathcal{X}}$. Then, each estimator $\widehat{\tau}_{h,\mathrm{tr}}^{s}$, $s\in\{10,00,11\}$, is (a) triply robust such that it is consistent for $\tau_{h}^{s}$ under $\mathcal{M}_{\pi+p}\cup\mathcal{M}_{\pi+\mu}\cup\mathcal{M}_{p+\mu}$; (b) asymptotically linear with influence function $\varphi_{h}^{s}$ under $\mathcal{M}_{\pi+p+\mu}$ such that it achieves the semiparametric efficiency lower bound.
By Theorem 2, $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. This triple robustness property of $\widehat{\tau}_{h,\mathrm{tr}}^{s}$ is obtained by separately investigating the properties of its numerator and the denominator. To account for the uncertainty in estimating the parametric nuisance functions, we estimate the variances of the triply robust estimators via the nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
We study further improvements to the parametric multiply robust estimator via debiased machine learning, which is similarly based on orthogonal moment functions (Chernozhukov2018) or EIFs, but allows efficient estimation of causal effects using data-adaptive nuisance estimators. Because the Neyman orthogonality holds for the EIF (Chernozhukov2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $n^{1/2}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates. Compared with parametric working models, machine learning based working models for the nuisance functions are more robust to potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting—is adopted. In this procedure, the entire sample is divided into separate parts for estimating the nuisance functions and the causal estimand, respectively (Chernozhukov2018). Mild regulatory conditions, mainly the consistency of machine learning first-step estimates, are required for the $n^{1/2}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (kennedy2023semiparametric).
We adopt the sample-splitting scheme for U-statistics provided in escanciano2022machine. Specifically, the sample index set $\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition $\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq n\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov2018 suggests setting $K$ to moderate values, such as 4 or 5, so that the sizes of elements in $\mathcal{I}$ are $O(n)$. We use the following example to demonstrate the definition of $\mathcal{I}$ and $\mathcal{P}$.
@@
With a slight abuse of notation, let $\widehat{\theta}=\{\widehat{\pi},\widehat{p}_{z},\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\widehat{\vartheta}=\{\widehat{\pi},\widehat{p}_{z}\}$ denote machine learning estimators of nuisance functions $\theta=\{\pi,p_{z},\mu_{z_{1}d_{1}z_{2}d_{2}}\}$ and $\vartheta=\{\pi,p_{z}\}$, respectively, as well as true functions $\theta^{*}=\{\pi^{*},p_{z}^{*},\mu_{z_{1}d_{1}z_{2}d_{2}}^{*}\}$ and $\vartheta^{*}=\{\pi^{*},p_{z}^{*}\}$. For a given partition $\mathcal{P}$ (based on the initial partition of singletons $\mathcal{I}$), we obtain intermediate estimates of the target estimand using observations in each $P_{l}$ (or $I_{k}$) with $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$) estimated using observations in $P_{l}^{c}$ (or $I_{k}^{c}$), those are not present in $\widehat{\theta}_{l}$ (or $\widehat{\vartheta}_{k}$). By pooling these intermediate estimates together, we obtain the debiased machine learning estimator for $\tau_{h}^{10}$ as
@@
The debiased machine learning estimators $\widehat{\tau}_{h,\mathrm{dml}}^{00}$ and $\widehat{\tau}_{h,\mathrm{dml}}^{11}$ can be constructed in a similar fashion and are given in Section S5 of the Online Supplement. To establish the asymptotic properties of $\widehat{\tau}_{h,\mathrm{dml}}^{s}$, we first assume $\theta$ and $\vartheta$ take values in $\Theta$, a convex subset of some normed vector space, and define $\Theta^{*}\subset\Theta$ as a nuisance realization set such that, for all $l$ and $k$, $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$ take values in this set with high probability. In addition, we assume true nuisance realizations $\theta^{*},\vartheta^{*}\in\Theta^{*}$. The following Theorem 3 summarizes the convergence and efficiency properties of the debiased machine learning estimators.
@@
    $g_{h,1}^{s}(\mathcal{O}_{1};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O}_{1},\mathcal{O}_{2};\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
then $\widehat{\tau}_{h,\mathrm{dml}}^{s}$ is $n^{1/2}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the variance of the EIF.
Theorem 3 establishes the consistency and asymptotic normality of the debiased machine learning estimators, building on the results developed for orthogonal moment functions given in Chernozhukov2022 and escanciano2022machine. Regulatory conditions in Theorem 3 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent, as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for existing machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator; similar conditions have been previously invoked in the debiased machine learning literature (Vansteelandt2022; Jiang2022). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (14) via the nonparametric bootstrap.